# The infrared luminosity of retired and post-starburst galaxies: A cautionary tale for star formation rate measurements

### v1

@@ Abstract
In galaxies with significant ongoing star formation there is an impressively tight correlation between total infrared luminosity ($L_{\mathrm{TIR}}$) and H $\alpha$ luminosity ($L_{\mathrm{H}\,\alpha}$), when H $\alpha$ is properly corrected for stellar absorption and dust attenuation. This long-standing result gives confidence that both measurements provide accurate estimates of a galaxy’s star formation rate (SFR), despite their differing origins. To test the extent to which this holds in galaxies with lower specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$, where $M_{*}$ is the stellar mass), we combine optical spectroscopy from the Sloan Digital Sky Survey (SDSS) with multi-wavelength (FUV to FIR) photometric observations from the Galaxy And Mass Assembly survey (GAMA). We find that $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ increases steadily with decreasing H $\alpha$ equivalent width ($W_{\mathrm{H}\,\alpha}$, a proxy for sSFR), indicating that both luminosities cannot provide a valid measurement of SFR in galaxies below the canonical star-forming sequence. For both ‘retired galaxies’ and ‘post-starburst galaxies’, $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ can be up to a factor of 30 larger than for star-forming galaxies. The smooth change in $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$, irrespective of star formation history, ionisation or heating source, dust temperature or other properties, suggests that the value of $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ is given by the balance between star-forming regions and ambient interstellar medium contributing to both $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$. While $L_{\mathrm{H}\,\alpha}$ can only be used to estimate the SFR for galaxies with $W_{\mathrm{H}\,\alpha}>3$ ($\mathrm{sSFR}\gtrsim 10^{-11.5}$/yr), we argue that the mid- and far-infrared can only be used to estimate the SFR of galaxies on the star-forming sequence, and in particular only for galaxies with $W_{\mathrm{H}\,\alpha}>10$ ($\mathrm{sSFR}\gtrsim 10^{-10.5}$/yr). We find no evidence for dust obscured star-formation in post-starburst galaxies.
@@ 1 Introduction
The basic physics of Balmer emission in star-forming galaxies is simple and well understood, with only photons with energies greater than 13.6 eV (Lyman continuum photons) able to contribute to the ionisation of the hydrogen atoms in the first place, meaning O and B stars are the primary contributor to Balmer emission in star-forming galaxies. However, there are two key complications in converting a H $\alpha$ emission line luminosity into a SFR. These are the attenuation of the photons as they travel through the galaxy ISM once leaving the H ii regions, and the contribution to the far-UV interstellar radiation field (ISRF) from active galactic nuclei (AGN), shocks and hot low-mass evolved stars (HOLMES; Flores-Fajardo et al., 2011), such as hot post-asymptotic giant branch stars and white dwarfs (Stasińska et al., 2008; Cid Fernandes et al., 2011). Through modelling the stellar continuum in SDSS galaxies using population synthesis models, Cid Fernandes et al. (2011) showed that galaxies with a low equivalent width of the H $\alpha$ line ($W_{\mathrm{H}\,\alpha}<3$) had an ionisation source consistent with the HOLMES present in the galaxy continuum spectrum, calling them ‘retired galaxies’. In these galaxies $L_{\mathrm{H}\,\alpha}$ is not proportional to their SFR, but is proportional to the total stellar mass of lower mass stars available to provide the ionising photons, as well as the presence of a reservoir of warm gas (see Herpich et al., 2018, for a comparison of retired galaxies with and without H $\alpha$ emission).
@@
On the other hand, the physics of mid- to far-IR luminosity is highly complex, with dependence on the gas density, dust-star geometry, the distribution of dust grain sizes and types which can vary with metallicity, and the shape of the ISRF, which depends on the stellar population balance (e.g. Draine & Li, 2007; Paladini et al., 2007; Bernard et al., 2010; Nersesian et al., 2019). It is predominantly the strong correlation with other measurements that provides the evidence for the link between TIR and SFR in high SFR galaxies. While Balmer emission lines are somewhat impacted by HOLMES, significant dust heating is caused by all stars with masses below O and B type stars. Diffuse dust emission permeates the entire galaxy contributing significantly to the IR luminosity of galaxies or regions with lower ongoing SFRs (e.g. Helou, 1986; Lonsdale Persson & Helou, 1987; Paladini et al., 2007; Bendo et al., 2010; Kennicutt et al., 2011; Calzetti, 2013; Nersesian et al., 2019). As a result, the TIR to H $\alpha$ luminosity ratio ($L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$) has been observed to increase with decreasing specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$, where $M_{*}$ is a galaxy’s stellar mass) in integrated SDSS–Herschel comparisons (Rosario et al., 2016) and spatially resolved PHANGS–MUSE comparisons (Belfiore et al., 2023). While Rosario et al. (2016) put the effect down to selection bias, Belfiore et al. (2023) interpreted it as due to contamination of the IR luminosity by diffuse ISM dust.
@@
As the debate about the role of total IR emission as a SFR indicator in quenching galaxies continues, and the origin of IR emission in ‘retired’ galaxies has not yet been studied, this paper takes a fresh look at the relationship between total IR luminosity and H $\alpha$ luminosity in a wide range of galaxies, including star-forming, post-starburst and retired galaxies. In Section 2 we present the catalogues that we have combined in our analysis, in Section 3 we summarise the relevant methods, and in Section 4 we present our results. We discuss our results with the help of a simple toy model in Section 5 and summarise in Section 6. Throughout the paper we assume a cosmology of $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$ and $h=0.7$ when converting fluxes into luminosity. All derived catalogues that we use assume a Chabrier (2003) initial mass function for conversion of stellar light into stellar masses. Unless otherwise stated $L_{\mathrm{H}\,\alpha}$ refers to dust attenuation and aperture corrected H $\alpha$ luminosity throughout the paper.
2 Data and sample selection
@@
From SDSS we take nebular emission line fluxes, errors and equivalent widths from the starlight–SDSS catalogue (Cid Fernandes et al., 2009; see also Mateus et al., 2006; Stasińska et al., 2006; Asari et al., 2007; Cid Fernandes et al., 2010). The underlying stellar continuum is subtracted prior to measurement of the emission line fluxes using the starlight full spectral fitting package (Cid Fernandes et al., 2005). Stellar masses are calculated from the mass-to-light ratio of the starlight model, and corrected for aperture loss using the fiber-to-total $z$-band luminosity fraction. The luminosity of the H $\alpha$ emission line is corrected for dust attenuation using the Cardelli et al. (1989) law with $R_{V}=3.1$ and assuming an intrinsic Balmer decrement of $\mathrm{H}\,\alpha/\mathrm{H}\,\beta=2.86$.
@@
We note that the GAMA magphys catalogue contains dust emission parameters even for galaxies with no mid- or far-IR photometric detections; these are predictions based on the magphys model priors. We therefore restrict the sample to only those galaxies with $\geq 2$ photometric bands between 22–500 µm detected at $>3\sigma$ in the lambdar catalogue. This reduces the sample to 4,368 galaxies. Finally, we restrict our sample to objects with H $\alpha$ detected at $>5\sigma$, leading to a final sample of 4,121 objects.
2.1 Post-starburst sample
@@
We match our combined SDSS+GAMA+H-ATLAS/WISE sample onto several SDSS DR7 post-starburst samples in the literature, identifying 33 post-starburst galaxies in total. Positions, catalogue identifiers and basic properties for these galaxies are given in Table 1. The sample includes 3 galaxies from Goto (2007), 24 from Wild et al. (2007) [^4] We follow Pawlik et al. (2018) and exclude dusty PCA-selected post-starburst candidates with H $\alpha$/H $\beta$ ratios $>6.6$. These are possible contaminants for this particular method, due to the impact of significant dust attenuation on the shape of the stellar continuum., 4 from Alatalo et al. (2016), 7 from Pattarakijwanich et al. (2016) and 9 applying the method described in Chen et al. (2019) to SDSS DR7, referred to as the ‘Tremonti sample’. Some objects are identified using multiple methods, with the largest cross over being 8 objects contained in both the Wild et al. (2007) and Tremonti samples. The selection criteria of each of these samples is summarised in Table 2.
@@
The sample includes 7 post-starburst galaxies where optical emission lines indicate that the dominant ionising source is residual star-formation (below the demarcation line of Stasińska et al. 2006), 8 are consistent with being ionised by HOLMES ($W_{\mathrm{H}\,\alpha}<3$), 6 ionised by AGN (above the demarcation line of Kewley et al. 2001 and $W_{\mathrm{H}\,\alpha}>3$), and 10 lie between the Stasińska et al. (2006) and Kewley et al. (2001) demarcation lines of the [N ii]/H $\alpha$ and [O iii]/H $\beta$ line ratios diagram (Baldwin et al., 1981). For the purposes of this paper it is essential to include galaxies with all ionisation sources, we therefore refrain from converting $L_{\mathrm{H}\,\alpha}$ into SFR for our figures, as in many cases this would be incorrect.
@@ 4 Results
The left panel of Figure 1 presents the total dust luminosity ($L_{\mathrm{TIR}}$), as estimated by magphys between 3–1000 µm versus the dust and aperture corrected H $\alpha$ luminosity ($L_{\mathrm{H}\,\alpha}$) for all galaxies in our sample, colour coded by the equivalent width of H $\alpha$ ($W_{\mathrm{H}\,\alpha}$). We immediately see that the expected correlation between $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ only holds for galaxies with $W_{\mathrm{H}\,\alpha}\gtrsim 3$Å, the limit below which a galaxy is classified as ‘retired’ (Cid Fernandes et al., 2011). The arrow indicates the direction of partial correlation ($\theta_{p}=-56^{\circ}$), confirming $W_{\mathrm{H}\,\alpha}$ to be strongly positively correlated with $L_{\mathrm{H}\,\alpha}$, and negatively with $L_{\mathrm{TIR}}$. As we will see from the next plot, the stronger correlation with $L_{\mathrm{H}\,\alpha}$ arises purely due to the presence of the retired galaxies.
@@
The right hand panel of Figure 1 restricts the sample to the 1,638 galaxies with optical emission line ratios that indicate the gas is predominantly ionised by young stellar populations, using the criteria of Stasińska et al. (2006) which is based on the commonly used [N ii]/H $\alpha$ and [O iii]/H $\beta$ line ratios. It is noteworthy that there is still a trend with $W_{\mathrm{H}\,\alpha}$, with the conversion line equating both luminosities to SFR only holding for galaxies with $W_{\mathrm{H}\,\alpha}\gtrsim 10$ (Hao et al., 2011; Murphy et al., 2011; Kennicutt & Evans, 2012). The partial correlation angle of $\theta_{p}=-48^{\circ}$ indicates that $W_{\mathrm{H}\,\alpha}$ traces $L_{\mathrm{TIR}}$ just as much as it does $L_{\mathrm{H}\,\alpha}$ and we can see that accounting for $W_{\mathrm{H}\,\alpha}$ would account for most of the scatter in the relation. The results shown in these figures are unchanged if we use a simple modified black body fit to the mid- and far-infrared data (see Section 3).
@@
In Figure 1 the post-starburst galaxies are indicated by larger circles. The identical colour scaling of the post-starburst galaxies relative to the background population of galaxies demonstrates that post-starburst galaxies are not unusual in their dust luminosities, given their H $\alpha$ equivalent widths. The majority of the post-starburst galaxies in this sample have $W_{\mathrm{H}\,\alpha}>3$, and in Figure 1 they lie on the upper side of the star-forming sequence in $L_{\mathrm{TIR}}$ versus $L_{\mathrm{H}\,\alpha}$. These are interpreted as being in the process of shutting down their star formation, or have alternative ionisation sources. A small number have $W_{\mathrm{H}\,\alpha}\lesssim 3$, i.e. consistent with being entirely quenched, and the majority of the line emission arising from older stars present in the galaxy. These are the objects with the highest excess $L_{\mathrm{TIR}}$.
@@
Our results show that all post-starburst galaxies have dust luminosities larger than predicted by $L_{\mathrm{H}\,\alpha}$, but consistent with star-forming galaxies or retired galaxies with the same $W_{\mathrm{H}\,\alpha}$. If these post-starburst galaxies are in fact ‘dust obscured starbursts’, then the excess $L_{\mathrm{TIR}}$ common to all galaxies with $W_{\mathrm{H}\,\alpha}<10$ should be interpreted as due to complete obscuration of the H $\alpha$ line. In the following sections we will discuss why we do not think this is the case.
@@
To further investigate the cause of the varying $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio with $W_{\mathrm{H}\,\alpha}$, Figure 2 shows the same samples plotted as stellar mass $M_{*}$ versus $L_{\mathrm{H}\,\alpha}$. The top left panel largely repeats the information presented in Figure 1, but the inclusion of stellar mass makes the blue and red sequences stand out more clearly. Interestingly, the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio changes smoothly across this figure, with no obvious discontinuity between the blue and red sequences, or for the post-starburst galaxies. The partial correlation angle of $\theta_{p}=-38^{\circ}$ is again consistent with $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ trending most closely with sSFR. The smoothness of the transition suggests that whatever is responsible for the varying ratio has no knowledge of the origin of the emission, or whether the galaxy is ‘quenched’, recently quenched or still star-forming.
Red sequence galaxies clearly have the highest $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$, with the majority having a ratio more than 1 dex greater than the expected ratio based on standard conversions (2.14, see Section 3), while 19 galaxies have a ratio $\geq 1.5$ dex greater. At the other extreme, galaxies with either very high $L_{\mathrm{H}\,\alpha}$/$M_{*}$ (equivalent to sSFR for galaxies where $L_{\mathrm{H}\,\alpha}$ is dominated by star-formation) or very high $L_{\mathrm{H}\,\alpha}$ have a ratio that is lower than expected, with 32 having a ratio $\geq 0.5$ dex lower. Again we see that post-starburst galaxies, whether completely quenched or not, have the expected $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio compared to other galaxies with the same stellar mass and $L_{\mathrm{H}\,\alpha}$.

### v2

@@ Abstract
In galaxies with significant ongoing star formation there is an impressively tight correlation between total infrared luminosity ($L_{\mathrm{TIR}}$) and H $\alpha$ luminosity ($L_{\mathrm{H}\,\alpha}$), when H $\alpha$ is properly corrected for stellar absorption and dust attenuation. This long-standing result gives confidence that both measurements provide accurate estimates of a galaxy’s star formation rate (SFR), despite their differing origins. To test the extent to which this holds in galaxies with lower specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$, where $M_{*}$ is the stellar mass), we combine optical spectroscopy from the Sloan Digital Sky Survey (SDSS) with multi-wavelength (FUV to FIR) photometric observations from the Galaxy And Mass Assembly survey (GAMA). We find that $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ increases steadily with decreasing H $\alpha$ equivalent width ($W_{\mathrm{H}\,\alpha}$, a proxy for sSFR), indicating that both luminosities cannot provide a valid measurement of SFR in galaxies below the canonical star-forming sequence. For both ‘retired galaxies’ and ‘post-starburst galaxies’, $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ can be up to a factor of 30 larger than for star-forming galaxies. The smooth change in $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$, irrespective of star formation history, ionisation or heating source, dust temperature or other properties, suggests that the value of $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ is determined by the balance between star-forming regions and ambient interstellar medium contributing to both $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$. It is not a result of the differing timescales of star formation that these luminosities probe. While $L_{\mathrm{H}\,\alpha}$ can only be used to estimate the SFR for galaxies with $W_{\mathrm{H}\,\alpha}>3$ ($\mathrm{sSFR}\gtrsim 10^{-11.5}$/yr), we argue that the mid- and far-infrared can only be used to estimate the SFR of galaxies on the star-forming sequence, and in particular only for galaxies with $W_{\mathrm{H}\,\alpha}>10$ ($\mathrm{sSFR}\gtrsim 10^{-10.5}$/yr). We find no evidence for dust obscured star-formation in local post-starburst galaxies.
@@ 1 Introduction
The basic physics of Balmer emission in star-forming galaxies is simple and well understood, with only photons with energies greater than 13.6 eV (Lyman continuum photons) able to contribute to the ionisation of the hydrogen atoms in the first place, meaning O and B stars are the primary contributor to Balmer emission in star-forming galaxies. However, there are two key complications in converting a H $\alpha$ emission line luminosity into a SFR. These are the attenuation of the photons as they travel through the galaxy ISM once leaving the H ii regions, and the contribution to the far-UV interstellar radiation field (ISRF) from active galactic nuclei (AGN), shocks and hot low-mass evolved stars (HOLMES; Flores-Fajardo et al., 2011), such as hot post-asymptotic giant branch stars and white dwarfs (Stasińska et al., 2008; Cid Fernandes et al., 2011). Through modelling the stellar continuum in SDSS galaxies using population synthesis models, Cid Fernandes et al. (2011) showed that galaxies with a low equivalent width of the H $\alpha$ emission line ($W_{\mathrm{H}\,\alpha}<3$) had an ionisation source consistent with the HOLMES present in the galaxy continuum spectrum, calling them ‘retired galaxies’. In these galaxies $L_{\mathrm{H}\,\alpha}$ is not proportional to their SFR, but is proportional to the total stellar mass of lower mass stars available to provide the ionising photons, as well as the presence of a reservoir of warm gas (see Herpich et al., 2018, for a comparison of retired galaxies with and without H $\alpha$ emission).
@@
On the other hand, the physics of mid- to far-IR luminosity is highly complex, with dependence on the gas density, dust-star geometry, the distribution of dust grain sizes and types which can vary with metallicity, and the shape of the ISRF, which depends on the stellar population balance (e.g. Draine & Li, 2007; Paladini et al., 2007; Bernard et al., 2010; Nersesian et al., 2019). It is predominantly the strong correlation with other measurements that provides the evidence for the link between TIR and SFR in high SFR galaxies. While Balmer emission lines are somewhat impacted by HOLMES, significant dust heating is caused by all stars with masses below O and B type stars. Diffuse dust emission permeates the entire galaxy contributing significantly to the IR luminosity of galaxies or regions with lower ongoing SFRs (e.g. Helou, 1986; Lonsdale Persson & Helou, 1987; Paladini et al., 2007; Bendo et al., 2010; Kennicutt et al., 2011; Calzetti, 2013; Nersesian et al., 2019). As a result, the TIR to H $\alpha$ luminosity ratio ($L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$) has been observed to increase with decreasing specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$, where $M_{*}$ is a galaxy’s stellar mass) in integrated SDSS–Herschel (Rosario et al., 2016) and SDSS–WISE (Salim et al., 2016) comparisons, and spatially resolved PHANGS–MUSE comparisons (Belfiore et al., 2023). While Rosario et al. (2016) put the effect down to selection bias, Belfiore et al. (2023) interpreted it as due to contamination of the IR luminosity by diffuse ISM dust.
@@
As the debate about the role of total IR emission as a SFR indicator in quenching galaxies continues, and the origin of IR emission in ‘retired’ galaxies has not yet been studied, this paper takes a fresh look at the relationship between total IR luminosity and H $\alpha$ luminosity in a wide range of galaxies, including star-forming, post-starburst and retired galaxies. In Section 2 we present the catalogues that we have combined in our analysis, in Section 3 we summarise the relevant methods, and in Section 4 we present our results. We discuss our results with the help of a simple toy model in Section 5 and summarise in Section 6. Throughout the paper we assume a cosmology of $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$ and $h=0.7$ when converting fluxes into luminosity. All derived catalogues that we use assume a Chabrier (2003) initial mass function for conversion of stellar light into stellar masses. Unless otherwise stated $L_{\mathrm{H}\,\alpha}$ refers to dust attenuation and aperture corrected H $\alpha$ luminosity.
2 Data
Figure 1: A summary of the properties of the galaxies studied in this paper. Galaxies are colour coded according to their emission line properties into star-forming (SF), AGN and retired galaxies, as described in the text. Post-starburst galaxies are identified via their stellar continuum properties, and are shown as darker coloured circles. Top: redshift vs. total stellar mass. Centre: [N ii]/H $\alpha$ vs.  [O iii]/H $\beta$ line flux ratios. Bottom: [N ii]/H $\alpha$ line flux ratio vs. the equivalent width of the H $\alpha$ emission line.
@@ 2 Data
From SDSS we take nebular emission line fluxes, errors and equivalent widths from the starlight–SDSS catalogue (Cid Fernandes et al., 2009; see also Mateus et al., 2006; Stasińska et al., 2006; Asari et al., 2007; Cid Fernandes et al., 2010). The underlying stellar continuum is subtracted prior to measurement of the emission line fluxes using the starlight full spectral fitting package (Cid Fernandes et al., 2005). Stellar masses are calculated from the mass-to-light ratio of the starlight model, and corrected for aperture loss using the fiber-to-total $z$-band luminosity fraction. The luminosity of the H $\alpha$ emission line is corrected for dust attenuation using the Cardelli et al. (1989) law with $R_{V}=3.1$ and assuming an intrinsic Balmer decrement of $\mathrm{H}\,\alpha/\mathrm{H}\,\beta=2.86$. This steeper dust law than e.g. Calzetti et al. (2000) is appropriate for correcting emission lines, due to the screen-like geometry of the dust in H ii regions (Wild et al., 2007; da Cunha et al., 2008; Wild et al., 2011). While the assumed dust law alters the absolute conversion between $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ in the case that both trace SFR, this does not impact the conclusions of this paper.
@@ 2.1 Sample selection
The GAMA magphys catalogue contains dust emission parameters even for galaxies with no mid- or far-IR photometric detections; these are predictions based on the magphys model priors. We therefore restrict our analysis to only those galaxies with $\geq 2$ photometric bands between 22–500 µm detected at $>3\sigma$ in the lambdar catalogue. This reduces the sample to 4,368 galaxies. Finally, we restrict our sample to objects with H $\alpha$ detected at $>5\sigma$, leading to a final sample of 4,121 galaxies with SDSS spectroscopy, GAMA photometry, H-ATLAS/WISE and H $\alpha$ emission line detections.
2.1.1 Star-forming and retired samples
Of these 4,121 galaxies, we find that 1,638 galaxies have optical emission line ratios that indicate the gas is solely ionised by young stellar populations, using the criteria of Stasińska et al. (2006), which is based on the commonly used [N ii]/H $\alpha$ and [O iii]/H $\beta$ line flux ratios. We find that 386 galaxies have H $\alpha$ equivalent widths $W_{\mathrm{H}\,\alpha}<3$Å, the limit below which a galaxy is classified as ‘retired’ i.e. the emission lines are consistent with the ionisation source being hot low-mass evolved stars (HOLMES) present in the galaxy continuum spectrum (Cid Fernandes et al., 2011). We also find 305 galaxies are classified as AGN using the Kewley et al. (2001) criteria. The ionising photons responsible for the H $\alpha$ emission in the remaining galaxies likely come from a mixture of HOLMES, ongoing star formation, shocks and low luminosity AGN. For the purposes of this paper it is essential to include galaxies with all ionisation sources, but we refrain from converting $L_{\mathrm{H}\,\alpha}$ into SFR for our figures. Figure 1 shows the samples in stellar mass vs. redshift, and on standard emission line diagnostic diagrams.
2.1.2 Post-starburst sample
@@
We match our combined SDSS+GAMA+H-ATLAS/WISE sample onto several SDSS DR7 post-starburst samples in the literature, identifying 33 post-starburst galaxies in total. Positions, catalogue identifiers and basic properties for these galaxies are given in Table 1. The sample includes 3 galaxies from Goto (2007), 24 from Wild et al. (2007) [^4] We follow Pawlik et al. (2018) and exclude dusty PCA-selected post-starburst candidates with H $\alpha$/H $\beta$ ratios $>6.6$. These are possible contaminants for this particular method, due to the impact of significant dust attenuation on the shape of the stellar continuum., 4 from Alatalo et al. (2016), 7 from Pattarakijwanich et al. (2016) and 9 applying the method described in Chen et al. (2019) to SDSS DR7, referred to as the ‘Tremonti sample’. Some objects are identified using multiple methods, with the largest cross over being 8 objects contained in both the Wild et al. (2007) and Tremonti samples. The selection criteria of each sample is summarised in Table 2.
@@
The sample includes 7 post-starburst galaxies where optical emission lines indicate that the dominant ionising source is residual star-formation (below the demarcation line of Stasińska et al. 2006), 8 are consistent with being ionised by HOLMES ($W_{\mathrm{H}\,\alpha}<3$), 6 ionised by AGN (above the demarcation line of Kewley et al. 2001 and $W_{\mathrm{H}\,\alpha}>3$), and 10 lie between the Stasińska et al. (2006) and Kewley et al. (2001) demarcation lines of the [N ii]/H $\alpha$ and [O iii]/H $\beta$ line ratios diagram (Baldwin et al., 1981). The post-starburst galaxies are shown with darker coloured symbols in Figure 1.
@@ 4 Results
The left panel of Figure 2 presents the total dust luminosity ($L_{\mathrm{TIR}}$), as estimated by magphys between 3–1000 µm, versus the dust and aperture corrected H $\alpha$ luminosity ($L_{\mathrm{H}\,\alpha}$) for all galaxies in our sample, colour coded by the equivalent width of H $\alpha$ ($W_{\mathrm{H}\,\alpha}$). We plot $W_{\mathrm{H}\,\alpha}$ as this is a direct observable; for galaxies with emission lines dominated by star formation $W_{\mathrm{H}\,\alpha}$ is strongly correlated with specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$). If both $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ were tracing SFR, galaxies should lie on the blue line. We can immediately see that this is only true for galaxies with $W_{\mathrm{H}\,\alpha}$ $\gtrsim 10$Å. Retired galaxies with $W_{\mathrm{H}\,\alpha}$ $\lesssim 3$Å (yellow to red points), lie significantly above or to the left the line, with higher $L_{\mathrm{TIR}}$ than expected for their $L_{\mathrm{H}\,\alpha}$. This can not be explained by contamination of the H $\alpha$ line by additional ionisation sources such as AGN, HOLMES or shocks, as such contamination would scatter galaxies to the right of the blue line. The arrow indicates the direction of partial correlation ($\theta_{p}=-55^{\circ}$), confirming $W_{\mathrm{H}\,\alpha}$ to be strongly positively correlated with $L_{\mathrm{H}\,\alpha}$, and negatively with $L_{\mathrm{TIR}}$.
@@
The right hand panel of Figure 2 restricts the sample to the 1,638 galaxies with optical emission line ratios that indicate the gas is predominantly ionised by young stellar populations and therefore $L_{\mathrm{H}\,\alpha}$ is a good proxy for SFR (Section 2.1.1). It is noteworthy that there is still a trend with $W_{\mathrm{H}\,\alpha}$, with the conversion line equating both luminosities to SFR only holding for galaxies with $W_{\mathrm{H}\,\alpha}\gtrsim 10$ (Hao et al., 2011; Murphy et al., 2011; Kennicutt & Evans, 2012). All other galaxies with lower $W_{\mathrm{H}\,\alpha}$ lie above the line, with higher $L_{\mathrm{TIR}}$ than expected for their $L_{\mathrm{H}\,\alpha}$. The partial correlation angle of $\theta_{p}=-48^{\circ}$ indicates that $W_{\mathrm{H}\,\alpha}$ traces $L_{\mathrm{TIR}}$ just as much as it does $L_{\mathrm{H}\,\alpha}$ and we can see that accounting for $W_{\mathrm{H}\,\alpha}$ would account for most of the scatter in the relation. The results shown in these figures are unchanged if we use a simple modified black body fit to the mid- and far-infrared data (see Section 3).
@@
In both panels of Figure 2 the post-starburst galaxies are indicated by larger circles. The identical colour scaling of the post-starburst galaxies relative to the background population of galaxies demonstrates that post-starburst galaxies are not unusual in their dust luminosities, given their H $\alpha$ equivalent widths. The majority of the post-starburst galaxies lie above the star-forming sequence in $L_{\mathrm{TIR}}$ versus $L_{\mathrm{H}\,\alpha}$. The small number with $W_{\mathrm{H}\,\alpha}\lesssim 3$ have the highest excess $L_{\mathrm{TIR}}$.
@@
Our results show that all post-starburst galaxies have dust luminosities larger than predicted by $L_{\mathrm{H}\,\alpha}$, but consistent with star-forming galaxies or retired galaxies with the same $W_{\mathrm{H}\,\alpha}$. If these post-starburst galaxies are in fact ‘dust obscured starbursts’, then the excess $L_{\mathrm{TIR}}$ common to all galaxies with $W_{\mathrm{H}\,\alpha}<10$ should be interpreted as due to complete obscuration of the H $\alpha$ line. In Section 5 we will discuss why we do not think this is the case.
@@
To further investigate the cause of the varying $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio with $W_{\mathrm{H}\,\alpha}$, Figure 3 shows the same samples plotted as stellar mass $M_{*}$ versus $L_{\mathrm{H}\,\alpha}$. We overplot lines of constant sSFR, although caution that these are only relevant where $L_{\mathrm{H}\,\alpha}$ arises predominantly from star formation. The top left panel largely repeats the information presented in Figure 2, but the inclusion of stellar mass makes the blue and red sequences stand out more clearly. Interestingly, the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio changes smoothly across this figure, with no obvious discontinuity between the blue and red sequences, or for the post-starburst galaxies. The partial correlation angle of $\theta_{p}=-38^{\circ}$ is consistent with $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ trending most closely with $L_{\mathrm{H}\,\alpha}$/$M_{*}$ (or sSFR for galaxies where $L_{\mathrm{H}\,\alpha}$ is dominated by star-formation). The smoothness of the transition suggests that whatever is responsible for the varying ratio has no knowledge of the origin of the emission, or whether the galaxy is quenched, recently quenched or still star-forming.
Red sequence galaxies clearly have the highest $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$, with the majority having a ratio more than 1 dex greater than the expected ratio based on standard conversions (2.14, see Section 3), while 19 galaxies have a ratio $\geq 1.5$ dex greater. If we accounted for contamination of the H $\alpha$ line from ionisation sources that are not related to star formation, such as AGN, shocks or HOLMES, this discrepancy would further increase. At the other extreme, galaxies with either very high $L_{\mathrm{H}\,\alpha}$/$M_{*}$ (high sSFR for galaxies where $L_{\mathrm{H}\,\alpha}$ is dominated by star-formation) or very high $L_{\mathrm{H}\,\alpha}$ have a ratio that is lower than expected, with 32 having a ratio $\geq 0.5$ dex lower. Again we see that post-starburst galaxies, whether completely quenched or not, have the expected $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio compared to other galaxies with the same stellar mass and $L_{\mathrm{H}\,\alpha}$.